# version 0.6

### v3

@@
position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments/suggestions welcome at nonstochastic.control@gmail.com
@@
The primary distinction between online nonstochastic control and other frameworks is the objective. In optimal control, robust control, and other control methodologies that assume stochastic noise, the goal is to perform comparably to an offline optimal strategy. In online nonstochastic control, both the cost functions as well as the perturbations from the assumed dynamical model are chosen by an adversary. Thus the optimal policy is not defined a priori. Rather, the target is to attain low regret against the best policy in hindsight from a benchmark class of policies.
This objective suggests the use of the decision making framework of online convex optimization as an algorithmic methodology. The resulting methods are based on iterative mathematical optimization algorithms and are accompanied by finite-time regret and computational complexity guarantees.
@@ 1.1 What is This Book About?
The advance of the deep learning revolution has brought to light the power and robustness of algorithms that are based on gradient methods. Deep learning pioneer and Turing award winner, professor Yann Lecun, is attributed the following quote from a Barbados workshop on deep learning theory, circa Feb 2020,
@@
This is a deep and difficult question that we fall short of answering. We do, however, propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient based control methods.
@@ 1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century. One of its most critical components was introduced by the Scottish inventor James Watt: amongst the earliest examples of a control mechanism, the Centrifugal Governor. A major source of losses in efficiency of earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily result in engine failure. The centrifugal governor was created to solve this problem.
@@ 1.3 Formalization and Examples of a Control Problem
The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$ and control input $\mathbf{u}_{t}$. The subscript indicates the relevant quantities at time step $t$.
Given a dynamical system, there are several possible objectives we can formulate. A generic formulation appears below, which captures many common control formulations, notably the optimal control problem and the robust control problem, which will be addressed in much more detail later in this book. In this problem formulation, the goal is to minimize the long-term horizon cost given by a sequence of cost functions.
@@
Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}\times d_{u}}\to{\mathbb{R}}$. This is given by the following mathematical program
@@ 1.3.1 Example: Control of a Medical Ventilator
The goal of ventilator control is to regulate pressure according to some predetermined pattern. As a dynamical system, we can write $p_{t+1}=f(p_{t},u_{t},w_{t}),$ and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$. The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
In the figure below we see an illustration of a simple controller, called the PID controller, as it performs on a ventilator and lung simulator.
Figure 1.2: Performance of the PID controller on a mechanical ventilator
@@ 1.4 Simple Control Algorithms
The following methods are simplistic but still widely used. For example, PID algorithm is still used in the majority of medical ventilators today. In later sections we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
@@ 1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$. A bang-bang controller chooses the control input at time $t$ as
@@ 1.6 The Need for a New Theory
We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and miss-specifications, and give us finite time provable guarantees? This is the subject of online non-stochastic control!
@@ 1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, i.e. arbitrary or even adversarial perturbations, the optimal policy is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can completely be predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games – regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$, and then observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
@@
    First, the best policy in hindsight may be very complicated to describe and reason about. This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
@@
    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs. the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed, as we do later in the book.
@@ 1.7 Bibliographic Remarks
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers an extensive coverage of topics from the viewpoint of dynamic programming. The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory. The excellent text of [SL91] gives a thorough introduction to adaptive control theory.
@@ 2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into a patient’s lungs. This inflow determines the volume of the lungs $v_{t}$ via a simple linear relation,
@@
The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-u_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort. See Figure 1.2 for a schematic representation of the airflow when a lung is connected to a mechanical ventilator.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics. The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@ 2.1.3 Pendulum Swing-up
$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\sin\theta}{ml^{2}},$
@@
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@ 2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system. The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevator angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
@@
Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia
@@
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0\\ 1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
@@ 2.1.5 Epidemiological Models
The COVID-19 pandemic marked an increased public interest in policy discussions on the effectiveness of certain interventions as a means of controlling the spread of infection. Epidemiological models forecast the spread of an infectious disease via structural models and differential equations. We present the discrete-time analogue of the most basic of such models – the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the susceptible, infected, and recovered population sizes. Let $u_{t}$ be the number of vaccines administered at time $t$.
@@
Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia
@@ 2.2.1 Equilibrium of Dynamical Systems
Clearly 00 is an equilibrium point as ${A}0=0$ for all ${A}$. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null veector. Thus 00 is a globally attractive, stable equilibrium.
In full generality, determining if a system has an equilibirum is an intractable problem, as we show in this chapter.
@@ 2.2.2 Stabilizability and Controllability
A Dynamical System is stabilizable if and only if, for every initial state, there exists a sequence of controls that drives the deterministic-equivalent system to the zero state. That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=0$.
@@ 2.2.3 Quantitative Definitions of Stabilizability
A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if and only if the following holds. For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t})).$

### v4

# version 0.75

@@
position=current page.north, angle=0, nodeanchor=north, vshift=-5mm, opacity=0.5, scale=1, contents=Draft; comments and suggestions welcome at nonstochastic.control@gmail.com
@@
The primary distinction between online nonstochastic control and other frameworks is the objective. In optimal control, robust control, and other control methodologies that assume stochastic noise, the goal is to perform comparably to an offline optimal strategy. In online nonstochastic control, both the cost functions as well as the perturbations from the assumed dynamical model are chosen by an adversary. Thus, the optimal policy is not defined a priori. Rather, the goal is to attain low regret against the best policy in hindsight from a benchmark class of policies.
This objective suggests the use of the decision-making framework of online convex optimization as an algorithmic methodology. The resulting methods are based on iterative mathematical optimization algorithms and are accompanied by finite-time regret and computational complexity guarantees.
Acknowledgments
This book was a journey that started in Princeton at the Google DeepMind Lab in the summer of 2018. At that time Sham Kakade visited the group, which consisted of the authors as well as Naman Agarwal, Brian Bullins, Xinyi Chen, Cyril Zhang and Yi Zhang, henceforth the gang. The gang worked on various aspects of optimization, reinforcement learning, and dynamical systems. The first paper, in which online nonstochastic control was basically invented, started in room 401 in the computer science building at Princeton University. We remember and cherish the exact moment along with the mathematical writings in white chalk on that blackboard.
Ever since then, the gem of an idea turned out into a complete theory which is now mainstream in machine learning and control, with numerous applications and refinements. Notable events that contributed to this theory include the following. Max Simchowitz, who was a graduate student then, visited Princeton during 2019-2020, where he collaborated with the group, and the theory was extended to partial observed systems. Paula Gradu, who was then an undergraduate, and Edgar Minasyan, while a grad student, contributed to the study of time-varying and nonlinear dynamical systems. Anirudha Majumdar joined for several important projects, notably for iterative linear control. Udaya Ghai and Arushi Gupta studied extending the GPC to model-free RL.
Most recently, this theory has been used to design new architectures for sequence prediction. The theory of spectral filtering was extended by the aforementioned members of the group, and newer folks joined, including Daniel Suo and Annie Marsden.
This book was further refined through teaching. Elad Hazan gratefully acknowledges the numerous contributions and insights of the students of the course computational control theory delivered at Princeton during COVID on Zoom during 2020-2021. The text was further enhanced by students of theoretical machine learning taught at Princeton University in the fall of 2024-2025.
The following students were particularly helpful making insightful comments and in finding various typos and suggesting improvements: Gon Buzaglo, Anand Brahmbhatt, Sofiia Druchyna (and many more we wish to announce very soon).
Throughout the writing of this book, we engaged in extensive discussions with ChatGPT, an AI assistant. Although it does not claim authorship, its ability to refine mathematical reasoning and clarify complex topics has been a remarkable collaboration in human-AI interaction. In fact, it wrote this acknowledgment all by itself!
@@ Acknowledgments
$K_{t}$ stabilizing linear controller at time $t$
$K_{r}(A,B)$ Kalman controllability matrix of order $r$
$\hat{K}(A,C)$ Kalman observability matrix
@@ 1.1 What is This Book About?
The advance of the deep learning revolution has brought to light the power and robustness of algorithms that are based on gradient methods. The following quote is attributed to deep learning pioneer Yann Lecun during a Barbados workshop on deep learning theory circa Feb 2020,
@@
This is a deep and difficult question that we fall short of answering. We do, however, propose a different way of answering this question: we define the control problem from the point of view of an online player in a repeated game. This gives rise to a simple but powerful framework called nonstochastic control theory. The framework is a natural fit to apply techniques from online convex optimization, resulting in new, efficient, gradient-based control methods.
@@ 1.2 The Origins of Control
The industrial revolution was in many ways shaped by the steam engine that came to prominence in the 18th century. One of its most critical components was introduced by the Scottish inventor James Watt: the centrifugal governor, which is amongst the earliest examples of a control mechanism. A major source of losses in efficiency of earlier designs was the difficulty in regulating the engine. Too much or too little steam could easily result in engine failure. The centrifugal governor was created to solve this problem.
@@ 1.3 Formalization and Examples of a Control Problem
The function $g_{t}$ is a mapping from the state to the observed state, and the function $f_{t}$ is the transition function, given current state $\mathbf{x}_{t}$, control input $\mathbf{u}_{t}$ and perturbation $\mathbf{w}_{t}$. The subscript indicates the relevant quantities at time step $t$.
Given a dynamical system, there are several possible objectives we can formulate. We give a generic formulation below that captures many common control formulations, notably the optimal control problem and the robust control problem. These will be addressed in much more detail later in this book. In this problem formulation, the goal is to minimize the long-term horizon cost given by a sequence of cost functions.
@@
Given a dynamical system with starting state $(\mathbf{x}_{1},\mathbf{y}_{1})$, iteratively construct controls $u_{t}$ to minimize a sequence of loss functions $c_{t}(\mathbf{y}_{t},\mathbf{u}_{t})$, $c_{t}:{\mathbb{R}}^{d_{y}+d_{u}}\to{\mathbb{R}}$. This is given by the following mathematical program
@@ 1.3.1 Example: Control of a Medical Ventilator
The goal of ventilator control is to regulate pressure according to some predetermined pattern. As a dynamical system, we can write $v_{t+1}=f(v_{t},u_{t},w_{t})$, $p_{t}=g(v_{t})$, and we can define the cost function to be $c_{t}(p_{t},u_{t})=|p_{t}-p_{t}^{*}|^{2},$ where $p_{t}^{*}$ is the desired pressure at time $t$. The goal, then, is to choose a sequence of control signals $\{u_{t}\}_{t=1}^{T}$ to minimize the total cost over $T$ iterations.
In the figure below we see an illustration of a simple controller, that we will soon describe, called the PID controller, as it performs on a ventilator and lung simulator.
Figure 1.2: Performance of the PID controller on a mechanical ventilator, from [SZG+21].
@@ 1.4 Simple Control Algorithms
The following methods are simplistic but still widely used. For example, the PID algorithm is still used in the majority of medical ventilators today. In later sections we describe a modern optimization-based framework that yields more sophisticated algorithms that are capable of better performance.
@@ 1.4.1 The Bang-Bang Controller
Bang-bang control is one of the simplest control strategies. Consider a scalar dynamical system, where it is desirable to keep the state $x_{t}$ within a pre-specified dynamic range $[x_{\min},x_{\max}]$ (that can vary with time), and where the scalar control $u_{t}$ input must at all times be within the interval $[u_{\min},u_{\max}]$, that is assumed to contain zero for simplicity. A bang-bang controller chooses the control input at time $t$ as
@@ 1.6 The Need for a New Theory
We would like an adaptive control theory which allows us to attain the best of both worlds: an efficient and fast flight when the weather permits, and a careful drone maneuver when this is required by the conditions. Can we design a control theory that will allow us to take into account the specific instance perturbations and misspecifications, and give us finite time provable guarantees? This is the subject of online non-stochastic control!
@@ 1.6.1 Online Nonstochastic Control Theory
This book is concerned with nonstochastic noise. When dealing with non-stochastic, i.e. arbitrary or even adversarial perturbations, the optimal policy, i.e. a decision making rule mapping observations to control inputs, is not clear a priori. Rather, an optimal policy for the observed perturbations can be determined only in hindsight. This is a significant difference from optimal and robust control theory, where the mapping from states to actions can completely be predetermined based upon the noise model.
Since optimality is defined only in hindsight, the goal is not predetermined. We thus shift to a different performance metric borrowed from the world of learning in repeated games – regret. Specifically, we consider regret with respect to a reference class of policies. In online control, the controller iteratively chooses a control $\mathbf{u}_{t}$. Then the controller observes the next state of the system $\mathbf{x}_{t+1}$ and suffers a loss of $c_{t}(\mathbf{x}_{t},\mathbf{u}_{t})$ according to an adversarially chosen loss function. For simplicity, we consider the case of full state observation where $\mathbf{y}_{t}=\mathbf{x}_{t}$. Let $\Pi=\{\pi:\mathbf{x}\to\mathbf{u}\}$ be a class of policies. The regret of the controller with respect to $\Pi$ is defined as follows.
@@
    First, the best policy in hindsight out of all possible policies may be very complicated to describe and reason about. This argument was made by the mathematician Tyrell Rockafellar, see the bibliographic section for more details.
@@
    Secondly, it can be shown that it is impossible, in general, to obtain sublinear regret vs. the best policy in hindsight. Rather, a different performance metric called competitive ratio can be analyzed, as we discuss later in the book.
@@ 1.7 Bibliographic Remarks
There are many excellent textbooks on control theory, which we point to in later chapters. The book of [Ber07] offers an extensive coverage of topics from the viewpoint of dynamic programming. The texts [Ste94, ZDG96a], on the other hand, delve deep into results on linear control, the most well-understood branch of control theory. Rockafellar [Roc87] proposed the use of convex cost functions for modeling the cost in a linear dynamical system to incorporate constraints on the state and control, and suggested computational difficulties with this approach. The text [SL91] gives a thorough introduction to adaptive control theory.
@@ 2.1.1 Medical Ventilator
In the task of mechanical ventilation, the control algorithm dictates the amount of oxygen inflow $u_{t}$ (in $\mathrm{m}^{3}/s$) into a patient’s lungs. This inflow approximately determines the volume of the lungs $v_{t}$ via a simple linear relation,
@@
The objective of the control algorithm is to track a physician-prescribed ideal pressure trajectory $p^{*}_{t}$ as closely as possible, say as measured by in squared loss $c_{t}(p_{t},u_{t})=|p_{t}-p^{*}_{t}|^{2}+|u_{t}|^{2}$, while limiting the maximum instantaneous airflow to minimize patient discomfort. See Figure 1.2 for a schematic representation of the airflow when a lung is connected to a mechanical ventilator.
@@ 2.1.2 Double Integrator
The double integrator is a common example of a linear dynamical system. In its simplest form, it models a one-dimensional point object moving according to Newtonian mechanics on a one dimensional trajectory (i.e. the real number line). The state $\mathbf{x}_{t}$ is 2-dimensional, with the first coordinate indicating the position of a point object and the second denoting its velocity. The control algorithm decides the force $u_{t}\in{\mathbb{R}}$ that is applied to the object at each time step. The dynamics equation can be written as
@@
$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}x^{*}\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@ 2.1.3 Pendulum Swing-up
$\theta_{t+1}=\theta_{t}+\Delta\dot{\theta}_{t},\qquad\dot{\theta}_{t+1}=\dot{\theta}_{t}+\Delta\frac{u_{t}-mg\ \sin\theta_{t}}{ml^{2}},$
@@
$c(\mathbf{x}_{t},u_{t})=\left\|\mathbf{x}_{t}-\begin{bmatrix}\pi\\ 0\end{bmatrix}\right\|^{2}+u_{t}^{2}.$
@@ 2.1.4 Aircraft Dynamics
A final example of a dynamical system we consider is the longitudinal (vertical tilt) motion of an aircraft that is regulated by automated controllers. We present the approximate dynamics for the Boeing 747 aircraft, which is a linear dynamical system, from [Boy10]. The 2-dimensional control input $\mathbf{u}_{t}$ corresponds to the elevator angle and the thrust. The disturbances correspond to the velocity of the wind along the body axis and perpendicular to it.
@@
Figure 2.4: Longitudinal control in an aircraft. Source: Wikipedia.
@@
$c(\mathbf{x}_{t},\mathbf{u}_{t})=\left\|\begin{bmatrix}1&0&0&0\\ 0&-1&0&7.74\end{bmatrix}\mathbf{x}_{t}-\mathbf{x}_{t}^{*}\right\|^{2}+\left\|\begin{bmatrix}0&1\end{bmatrix}\mathbf{u}_{t}\right\|^{2}.$
@@ 2.1.5 Epidemiological Models
The COVID-19 pandemic marked an increased public interest in policy discussions on the effectiveness of certain interventions as a means of controlling the spread of infection. Epidemiological models forecast the spread of an infectious disease via structural models and differential equations. We present the discrete-time analogue of the most basic of such models – the SIR model. Let $S_{t}$, $I_{t}$, and $R_{t}$ denote the susceptible, infected, and recovered population sizes. Let $u_{t}$ be the number of vaccines administered at time $t$. The dynamics of the basic SIR model are given by
@@
Figure 2.5: Trajectory of a continuous-time SIR model. Source: Wikipedia.
@@ 2.2.1 Equilibrium of Dynamical Systems
Clearly $\mathbf{0}$ is an equilibrium point as ${A}\mathbf{0}=\mathbf{0}$ for all ${A}$. If the modulus of all eigenvalues is strictly less than one, then after sufficiently many applications of $A$ any vectors converges to the null vector. Thus $\mathbf{0}$ is a globally attractive, stable equilibrium.
In full generality, determining if a system has an equilibrium is an intractable problem, as we show in this chapter.
@@ 2.2.2 Stabilizability and Controllability
A Dynamical System is stabilizable if and only if, for every initial state, there exists a sequence of controls that drives the deterministic-equivalent system to the zero state. That is, $\lim_{t\rightarrow\infty}{\mathbf{x}_{t}}=\mathbf{0}$.
@@ 2.2.3 Quantitative Definitions of Stabilizability
A control policy $\pi$ for a dynamical system $f$ is $\gamma$-stabilizing if and only if the following holds. For every starting state $\|\mathbf{x}_{0}\|\leq 1$ and every sequence of adversarially chosen bounded perturbations $\mathbf{w}_{1:T}$ such that $\|\mathbf{w}_{t}\|\leq 1$ for all $t\in[T]$, let $\mathbf{x}_{t}^{\pi}$ be the state obtained by using policy $\pi$,
$\mathbf{x}_{t+1}^{\pi}=f(\mathbf{x}_{t}^{\pi},\pi(\mathbf{x}_{t}),\mathbf{w}_{t}).$